Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
Gherardini - 15:55 Wednesday 23 January 2019 (44495) Print this report
Operator report - morning shift 23/1
The morning I found the itf locked in LOW_NOISE_1 state; the main activity was to recover the DC readout lock, Michal, Romain and Romain worked all the morning trying to recover the alignment of the B1 photodiodes: first with the itf in LOW_NOISE_1 state then with only the NI reflection; at the end the opening of SDB2 minitower was needed, its venting started at around 12:30utc and the minitower opened, the activity on the SDB2 is still in progress...


-inf
at 13:00utc I set the detection lab UTA in "portata nominale" (SDB2 minitower opening)
Search
×

Warning

×